OpenStudy (anonymous):

Solve the following system of equations by substitution. 3y = 9 9y + 2x = -12 A. (1/3, 15/2) B. (-1/3, - 15/2) C. (8, 8) D. (-39/2, 3)

OpenStudy (anonymous):

\(y=9/3=3\)

OpenStudy (anonymous):

take first equation, solve for y. y=9/3=3 substitute in the second 9(9/3) + 2x = -12 solve for x ==> x= -39/2 so, it's D.

OpenStudy (anonymous):

Thank your guys so much for your help! :)

OpenStudy (anonymous):

Put \(y\) in the second equation and you get : \(2x+27=-12\) So: \(x=-39/2\) \((-39/2,3)\)